# HALC: Object Hallucination Reduction via Adaptive Focal-Contrast Decoding

### v1

@@ 1 Introduction
The confluence of natural language processing (NLP) and computer vision (CV) has undergone a transformative shift over the past years with the introduction of vision-language models (VLMs) (Long et al., 2022; Zhu et al., 2023; Liu et al., 2023b). Although VLMs have shown exceptional proficiency in integrating and interpreting intricate data across both textual and visual modalities, a significant challenge emerged as the phenomenon of object hallucination (OH), where VLMs erroneously generate hallucinated objects and descriptions within their outputs (Rohrbach et al., 2018). Based on the different parts of the sentences that are being hallucinated, OH can be categorized into three types: object existence, attribute, and relationship hallucinations (Gunjal et al., 2023; Zhai et al., 2023).
@@
To this end, we introduce Object Hallucination Reduction through Adaptive FocaL-Contrast decoding (HALC), a novel decoding strategy designed to effectively counter OH and can be easily integrated into any open-source LVLMs such as MiniGPT-4 (Chen et al., 2023), LLaVA (Liu et al., 2023b) and mPLUG-Owl2 (Ye et al., 2023). HALC addresses all three types of OH (existence, attribute, and relationship) while preserving linguistic quality in both local and global levels; locally, it employs an adaptive focal-contrast grounding mechanism to locate the fine-grained optimal visual information to correct each generated token that might be hallucinating; and globally, it incorporates a matching-based beam search that utilizes a visual matching score to steer the generation of the final outputs to balance both OH mitigation and text generation quality.
The main contributions of this paper are: (1) HALC, a novel, plug-and-play decoding algorithm that significantly reduces OH in LVLMs while preserving outputs generation quality; (2) an open-sourced platform that unifies all major OH reduction baselines and state-of-the-arts (SOTAs) (Chuang et al., 2023; Zhou et al., 2023; Yin et al., 2023; Huang et al., 2023; Leng et al., 2023), including HALC, into one framework providing convenient evaluations supporting major LVLM backbones (Zhu et al., 2023; Chen et al., 2023; Liu et al., 2023b; Dai et al., 2023) and OH benchmarks and evaluation metrics (Rohrbach et al., 2018; Fu et al., 2023; Li et al., 2023; Liu et al., 2023a) and (3) comprehensive experimental studies that thoroughly evaluates HALC, demonstrating its superior capability in OH reduction over existing approaches.
@@ 2 Related Work
OH and its assessment. OH refers to the phenomenon where vision-language models (VLMs), including both the earlier BERT-based models (Li et al., 2019; Radford et al., 2021) and the more recent LVLMs (Liu et al., 2023b; Zhu et al., 2023; Tu et al., 2023; Cui et al., 2023; Wang et al., 2024), erroneously generate unfaithful contents. More specifically, Gunjal et al. (2023) and Zhai et al. (2023) proposed that OH could be categorized into three types: object existence hallucination for the creation of non-existent objects, object attribute hallucination for providing misleading descriptions, and object relationship hallucination for depicting incorrect inter-object relationships.
@@
Challenges and existing approaches. OH has been a persistent challenge over the past years (Rohrbach et al., 2018). Despite numerous advancements in LVLMs (Dai et al., 2022; Li et al., 2023; Zhou et al., 2024), none of them can produce faithful outputs without suffering from some level of OH. Various strategies have been developed to this matter. For instance, Zhou et al. (2023) and Yin et al. (2023) proposed post-hoc and self-correction pipelines, respectively. Huang et al. (2023) and Leng et al. (2023) developed decoding strategies emphasizing better prior utilization. While effective, these approaches often require powerful external LVLMs or additional data, limiting their adaptability.
@@ 3.3 Fine-grained Visual Knowledge Reduces OH
We verify our hypothesis through an empirical pilot study. Fig. 1 shows the oracle performance of OH levels when we rely on optimal visual contexts for tokens through brute-force search, with greedy decoding on the MME benchmark (Fu et al., 2023) on three categories of OH. [^1] Details of this oracle analysis can be found in Appendix B.2 We can see that for most cases, there are optimal visual contexts where decoding from them eliminates over 84.5% of the hallucinations. This motivates our approach of identifying different visual contexts for object-related token generations through adaptive focal-contrast decoding, which is introduced in detail in the next section.
@@ 4.3 Adaptive Focal-contrast Grounding
FOV sampling. We first sample a sequence of $n$ FOVs, $v_{1},v_{2},\ldots,v_{n}$, based on the initial visual context $v_{d}$. There could be different approaches to come up with different FOVs conditioning on $v_{d}$. To attain a larger coverage of the input image quickly, one strategy to sample FOVs is through an exponential expanding function, by setting $v_{i}=(w_{i},h_{i},p_{i})=\left((1+\lambda)^{i}w_{d},(1+\lambda)^{i}h_{d},p_{d}\right)$, where $w_{i},h_{i},p_{i}$ are the width, height, and center of the FOV $v_{i}$.
@@ 4.4 Matching-based Beam Search
Combining all components, the full procedure of HALC is summarized in Algorithm 1. Notice that by utilizing the fine-grained visual information at different levels for a single generation step, we admittedly trade in some computation time for correcting token hallucinations. One way to increase the HALC decoding speed is through parallelization of decoding from different visual contexts, where we can hope to spend at worst roughly twice of the regular decoding time at HALC steps considering the whole sequence. [^4] As HALC does not happen at every decoding step. There are also other overhead such as visual grounding affecting the runtime.
@@
18:        end fornormal-$\triangleright$ $y_{\text{new}}^{b}$ with $2m$ candidates obtained
@@ 6 Experiments
Baselines. To effectively evaluate HALC, besides regular greedy decoding and beam search baselines, we further involve layer-wise contrastive decoding SOTA DoLa (Chuang et al., 2023), as well as SOTA methods specifically designed to mitigate OH, including OPERA (Huang et al., 2023), VCD (Leng et al., 2023), Woodpecker (Yin et al., 2023) and LURE (Zhou et al., 2023) in our analysis. All the results are acquired and benchmarked consistently with our unified implementation. Please refer to Appendix B.1 for the detailed setting of our HALC.
@@ 6.1 CHAIR and POPE on MSCOCO
We also adjust the main metrics for comparison. As it is more random for descriptions to include the exact sampled hallucinated objects, false-negative (FN) and the resulting recall become less trustable in the offline checks. Therefore, we propose to use F-beta, instead of F-1, as the main metric of OPOPE, so that the final score relies less on the FN. Specifically, we have $F_{\beta}=(1+\beta^{2})\cdot(\text{precision}\cdot\text{recall})/(\beta^{2}\cdot\text{precision}+\text{recall})$, where we use $\beta=0.2$ throughout our experiments. The evaluation results incorporating OPOPE is shown in Table 2. All the numbers are averaged results of the three sampling methods (random, popular and adversarial, as in the original POPE), while the complete version of the table is shown in Appendix D. HALC outperforms other methods in most of the settings.
@@ 6.2 MME
The Multimodal Large Language Model Evaluation (MME) (Fu et al., 2023) benchmark is a comprehensive tool designed to quantitatively compare multimodal LLMs. Following Yin et al. (2023); Leng et al. (2023), we utilize the “existence” and “count” subsets to evaluate the object existence hallucinations and the “position” and “color” subsets for object attribute and relationship hallucination. Please refer to Appendix C for experiment details. The comprehensive results across six methods are reported in Fig. 5, where HALC significantly outperforms all the other methods on each sub-task, indicating an overall performance gain in reducing OH while preserving generation quality.
@@ 6.3 LLaVA-Bench Qualitative Study
LLaVA-Bench (Liu et al., 2023a) is a collection of 24 images, where each image is paired with a detailed, manually-crafted description and carefully selected questions. The questions are divided into three categories: simple QA (conversation), detailed descriptions, and complex reasoning. In this experiment, we leverage LLaVA-Bench as a case study to qualitatively compare the decoding outputs of HALC with other methods. The results are shown in Appendix E.
@@ Acknowledgement
We thank Lingyu Gao for initial discussion and constructive suggestions. This work was supported in part by the Research Computing Center at the University of Chicago. We also thank Center for AI Safety and Google Cloud Research Credits program for supporting our computing needs. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of any funding agencies.
@@
$*\pi(\cdot|v_{d})\rightarrow v^{*}$: Thus we consider sampling conditioned on $v_{d}$ in the FOV space to enhance the robustness of optimal visual context approximation, hoping to find some sample that is close to the optimal. To do this, we obtain an upper bound on the minimum deviation from the output distribution among a collection of FOV samples. Assume $\pi(\cdot|v_{d})\in\Omega$ is an arbitrary sampling function conditional on the initial FOV detection $v_{d}$, where $\Omega$ denotes the sampling space over all potential visual contexts in the entire image $v$. $\pi$ can either be a deterministic sampling function, or a stochastic sampling process with a probabilistic distribution over $\Omega$. Suppose we acquire $n$ samples $v_{1},v_{2},\ldots,v_{n}$ according to $\pi(\cdot|v_{d})$, we denote the minimum deviation of the resulted token probability from that of the optimal visual context $v^{*}$ as
@@
Appendix B Experimentation Details
B.1 Experimental Setups
@@
B.2 Empirical Studies on Optimal Visual Contexts
@@
Appendix C MME Experiment Details
The experiment details mostly follow Appendix B.2, where we adopt each sub-task of 30 images from the MME benchmark dataset  [^13] https://github.com/BradyFU/Awesome-Multimodal-Large-Language-Models/tree/Evaluation, and reconstruct the question prompt following offline POPE. Specifically, instead of simply asking a question with a binary yes/no answer, we first ask the decoder to generate a detailed caption of the provided image and then check whether the target positive/negative word existes in the caption. The detailed results are reported in Table 8. The corresponding figure result is shown in Fig. 5.

### v2

@@ 1 Introduction
The confluence of natural language processing (NLP) and computer vision (CV) has undergone a transformative shift over the past years with the introduction of vision-language models (VLMs) (Zhu et al., 2023; Liu et al., 2023b; Zhang et al., 2024). Although VLMs have shown exceptional proficiency in integrating and interpreting intricate data across both textual and visual modalities, a significant challenge emerged as the phenomenon of object hallucination (OH), where VLMs erroneously generate hallucinated objects and descriptions within their outputs (Rohrbach et al., 2018). Based on the different parts of the sentences that are being hallucinated, OH can be categorized into three types: object existence, attribute, and relationship hallucinations (Gunjal et al., 2023; Zhai et al., 2023).
@@
To this end, we introduce Object Hallucination Reduction through Adaptive FocaL-Contrast decoding (HALC), a novel decoding strategy designed to effectively counter OH and can be easily integrated into any open-source LVLMs such as MiniGPT-4 (Chen et al., 2023), LLaVA (Liu et al., 2023b) and mPLUG-Owl2 (Ye et al., 2023). HALC addresses all three types of OH (existence, attribute, and relationship) while preserving generation quality in both local and global levels; locally, it employs an adaptive focal-contrast grounding mechanism to locate the fine-grained optimal visual information to correct each generated token that might be hallucinating; and globally, it incorporates a matching-based beam search that utilizes a visual matching score to steer the generation of the final outputs to balance both OH mitigation and text generation quality.
The main contributions of this paper are: (1) HALC, a novel, plug-and-play decoding algorithm that significantly reduces OH in LVLMs while preserving outputs generation quality; (2) an open-sourced platform that unifies all major OH reduction baselines and state-of-the-arts (SOTAs) (Chuang et al., 2023; Zhou et al., 2023; Yin et al., 2023; Huang et al., 2023; Leng et al., 2023), including HALC, into one framework providing convenient evaluations supporting major LVLM backbones (Zhu et al., 2023; Chen et al., 2023; Liu et al., 2023b; Dai et al., 2023) and OH benchmarks and evaluation metrics (Rohrbach et al., 2018; Fu et al., 2023; Li et al., 2023; Liu et al., 2023a); and (3) comprehensive experimental studies that thoroughly evaluates HALC, demonstrating its superior capability in OH reduction over existing approaches.
@@ 2 Related Work
OH and its assessment. OH refers to the phenomenon where vision-language models (VLMs), including both earlier BERT-based models (Li et al., 2019; Radford et al., 2021) and more recent LVLMs (Liu et al., 2023b; Zhu et al., 2023; Tu et al., 2023; Cui et al., 2023; Wang et al., 2024; Zhou et al., 2024b), erroneously generate unfaithful contents. More specifically, Gunjal et al. (2023) and Zhai et al. (2023) proposed that OH could be categorized into three types: object existence hallucination for the creation of non-existent objects, object attribute hallucination for providing misleading descriptions, and object relationship hallucination for depicting incorrect inter-object relationships.
@@
Challenges and existing approaches. OH has been a persistent challenge over the past years (Rohrbach et al., 2018). Despite numerous advancements in LVLMs (Dai et al., 2022; Li et al., 2023; Zhou et al., 2024a), none of them can produce faithful outputs without suffering from some level of OH. Various strategies have been developed to this matter. For instance, Zhou et al. (2023) and Yin et al. (2023) proposed post-hoc and self-correction pipelines, respectively. Huang et al. (2023) and Leng et al. (2023) developed decoding strategies emphasizing better prior utilization. While effective, these approaches often require powerful external LVLMs or additional data, limiting their adaptability.
@@ 3.3 Fine-grained Visual Knowledge Reduces OH
We verify our hypothesis through an empirical pilot study. Fig. 1 shows the oracle performance of OH levels when we rely on optimal visual contexts for tokens through brute-force search, with greedy decoding on the MME benchmark (Fu et al., 2023) on three categories of OH. [^1] Details of this oracle analysis can be found in Appendix C.2. We can see that for most cases, there are optimal visual contexts where decoding from them eliminates over 84.5% of the hallucinations. This motivates our approach of identifying different visual contexts for object-related token generations through adaptive focal-contrast decoding, which is introduced in detail in the next section.
@@ 4.3 Adaptive Focal-contrast Grounding
FOV sampling. We first sample a sequence of $n$ FOVs, $v_{1},v_{2},\ldots,v_{n}$, based on the initial visual context $v_{d}$. There could be different approaches to come up with different FOVs conditioning on $v_{d}$. To attain a larger coverage of the input image quickly, one strategy to sample FOVs is through an exponential expanding function, by setting
$v_{i}=(w_{i},h_{i},p_{i})=\left((1+\lambda)^{i}w_{d},(1+\lambda)^{i}h_{d},p_{d}\right)$ (2)
where $w_{i},h_{i},p_{i}$ are the width, height, and center of FOV $v_{i}$.
@@ 4.4 Matching-based Beam Search
Combining all components, the full procedure of HALC is summarized in Algorithm 1. Notice that by utilizing the fine-grained visual information at different levels for a single generation step, we admittedly trade in some computation time for correcting token hallucinations. The detailed analysis on time complexity is in Appendix B. One way to increase the HALC decoding speed is through parallelization of decoding from different visual contexts, where we can hope to spend at worst roughly twice of the regular decoding time at HALC steps considering the whole sequence. [^4] As HALC does not happen at every decoding step. There are also other overhead such as visual grounding affecting the runtime.
@@
18:        end $\triangleright$ $y_{\text{new}}^{b}$ with $2m$ candidates obtained
@@ 6 Experiments
Baselines. To effectively evaluate HALC, besides regular greedy decoding and beam search baselines, we further involve layer-wise contrastive decoding SOTA DoLa (Chuang et al., 2023), as well as SOTA methods specifically designed to mitigate OH, including OPERA (Huang et al., 2023), VCD (Leng et al., 2023), Woodpecker (Yin et al., 2023) and LURE (Zhou et al., 2023) in our analysis. All the results are acquired and benchmarked consistently with our unified implementation. Please refer to Appendix C.1 for the detailed setting of our HALC.
@@ 6.1 CHAIR and POPE on MSCOCO
We also adjust the main metrics for comparison. As it is more random for descriptions to include the exact sampled hallucinated objects, false-negative (FN) and the resulting recall become less trustable in the offline checks. Therefore, we propose to use F-beta, instead of F-1, as the main metric of OPOPE, so that the final score relies less on the FN. Specifically, we have $F_{\beta}=(1+\beta^{2})\cdot(\text{precision}\cdot\text{recall})/(\beta^{2}\cdot\text{precision}+\text{recall})$, where we use $\beta=0.2$ throughout our experiments. The evaluation results incorporating OPOPE is shown in Table 2. All the numbers are averaged results of the three sampling methods (random, popular and adversarial, as in the original POPE), while the complete version of the table is shown in Appendix F. We also include the original POPE evaluation results in Appendix E where HALC also outperforms other methods in most of the settings.
@@ 6.2 MME
The Multimodal Large Language Model Evaluation (MME) (Fu et al., 2023) benchmark is a comprehensive tool designed to quantitatively compare multimodal LLMs. Following Yin et al. (2023); Leng et al. (2023), we utilize the “existence” and “count” subsets to evaluate the object existence hallucinations and the “position” and “color” subsets for object attribute and relationship hallucination. Please refer to Appendix D for experiment details. The comprehensive results across six methods are reported in Fig. 5, where HALC significantly outperforms all the other methods on each sub-task, indicating an overall performance gain in reducing OH while preserving generation quality.
@@ 6.3 LLaVA-Bench Qualitative Study
LLaVA-Bench (Liu et al., 2023a) is a collection of 24 images, where each image is paired with a detailed, manually-crafted description and carefully selected questions. The questions are divided into three categories: simple QA (conversation), detailed descriptions, and complex reasoning. In this experiment, we leverage LLaVA-Bench as a case study to qualitatively compare the decoding outputs of HALC with other methods. The results are shown in Appendix G.
7 Analysis and Ablation Studies
7.1 Adaptive Focal-contrast Grounding
FOV Sampling initialization. The visual context retrieval process described in §4.2 utilizes detector output as a key component of the adaptive focal-contrast grounding algorithm introduced in §4.3. However, it is important to note that HALC primarily uses the detector output as a initialization for the field of view (FOV) sampling process, rather than depending heavily on it. In this section, we present empirical results to compare different methods of sampling initialization, which include random sampling (selecting a random FOV within the image), center initialization (selecting a fixed region in the center of the image), original image initialization (using the entire image) and detector initialization (using the detector output). More specifically, we include an extra detector model, OWLv2 (Minderer et al., 2024), in addition to the Grounding Dino (Liu et al., 2023c) illustrated in previous sections.
Table 3: HALC performance with different sampling initialization.
[…]
Random 25.6 11.8 83.33 67.67 15.10
Center 23.9 11.2 86.62 69.10 14.80
Original 27.8 12.2 85.20 68.33 15.50
G. Dino 22.0 8.8 88.20 70.67 16.40
OWLv2 23.4 10.8 84.47 67.50 15.70
As shown in Table 3, both random and center initialization perform better than using the original image as the visual input. This result confirms the robustness of the proposed FOV sampling process. Additionally, both detectors deliver better performance than the other initializations, further demonstrating that using a detector-grounded FOV provides an effective starting point for the subsequent conditional FOV sampling process.
Exponential Expanding ratio. Besides initialization, another important parameter used in adaptive focal-contrast grounding is the expanding ratio $\lambda$, which determines each sampling FOV as in Eq. (2). Thus we further analyze the performance of HALC with different expanding ratios.
Table 4: HALC performance with different expanding ratios.
[…]
0.2 22.0 8.5 86.45 69.63 16.60
0.4 18.0 7.6 87.33 70.20 16.10
0.6 22.0 8.8 88.20 70.67 16.40
0.8 28.0 9.6 86.45 69.63 14.80
1.0 26.0 8.9 84.32 69.63 14.70
Table 4 demonstrates that an expanding ratio of 0.6 is optimal. We hypothesize that the poorer performance associated with smaller or larger expanding ratios is due to that smaller ratios increase the number of FOV samples, which presents greater challenges for the global beam search. On the other hand, larger ratios decrease the granularity of the FOV in the image, potentially leading to more severe hallucinations.
7.2 Global Beam Search
Beam sizes. As is common with all beam search algorithms, beam size $k$ is a major hyperparameter. Thus here we examine the performance of HALC w.r.t. different values of $k$.
Table 5: HALC performance with different values of beam size $k$.
[…]
1 36.0 14.6 88.20 70.49 15.40
2 22.0 8.8 88.74 70.67 16.40
3 26.0 9.8 87.65 70.67 15.40
5 29.6 11.1 86.33 70.14 15.70
8 33.3 13.8 87.73 70.14 15.50
Table 5 shows improved performance as the beam size initially increases from one. However, when the beam size reaches or exceeds two, the number of FOV samples also increases, making it more challenging for the global beam search module to select the optimal visual context from all the samples, thus leading to a higher rate of hallucination. Furthermore, as the beam size continues to increase, the variance of HALC’s performance also increases, indicating that it will be more difficult to select the top candidate as the global matching model also suffers from hallucination.
Scoring methods. Finally, we compare the BLIP and CLIP scoring models with random selection to rank the beams.
Table 6: HALC performance with different scoring methods.
[…]
Random 26.6 12.8 85.45 68.45 15.20
BLIP 22.0 8.8 88.20 70.67 16.40
CLIP 23.4 10.0 87.67 71.96 15.60
As shown in Table 6, different scoring methods do not lead to large variations and they all outperform random selection.
@@ 8 Conclusion
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Acknowledgement
We thank Lingyu Gao for initial discussion and constructive suggestions. This work was supported in part by the Research Computing Center at the University of Chicago, and Cisco Faculty Research Award. We also thank Center for AI Safety and Google Cloud Research Credits program for supporting our computing needs. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of any funding agencies.
@@
$\pi(\cdot|v_{d})\rightarrow v^{*}$: Thus we consider sampling conditioned on $v_{d}$ in the FOV space to enhance the robustness of optimal visual context approximation, hoping to find some sample that is close to the optimal. To do this, we obtain an upper bound on the minimum deviation from the output distribution among a collection of FOV samples. Assume $\pi(\cdot|v_{d})\in\Omega$ is an arbitrary sampling function conditional on the initial FOV detection $v_{d}$, where $\Omega$ denotes the sampling space over all potential visual contexts in the entire image $v$. $\pi$ can either be a deterministic sampling function, or a stochastic sampling process with a probabilistic distribution over $\Omega$. Suppose we acquire $n$ samples $v_{1},v_{2},\ldots,v_{n}$ according to $\pi(\cdot|v_{d})$, we denote the minimum deviation of the resulted token probability from that of the optimal visual context $v^{*}$ as
@@
Appendix B Time Cost Analysis
Since time complexity is a critical aspect for VLM decoding algorithms, in this section we analyze the additional runtime overhead and time cost of HALC. According to Biber et al. (2000), nouns, adjectives, adverbs, numbers, verbs, and pronouns, which are tokens that will actually pass through HALC decoding, comprise approximately 35% of the total words in modern English (we observe similar sparse patterns in our experiments). POS tagging is observably fast in practice (we used the spaCy package, which is highly optimized on CPU with the smallest tagger model, which is only 12 MB in size [^7] https://spacy.io/models/en#en_core_web_sm). Thus we will mainly discuss the time cost w.r.t. other modules in HALC.
For each individual token, after its original decoding, HALC will utilize the detection module to initialize the FOV sampling, for which we use $T_{d}$ to represent the detector time cost. Next, each one of the $n$ FOVs (in our experiments, $n=4$, as shown in Table 8) is fed back into the LVLM for decoding, resulting in $n*T_{LVLM}$ time cost, where $T_{LVLM}$ represents the LVLM decoding time for a single step (although this may increase slightly as the sequence grows longer). Other computations on top of the multiple decodings such as contrasting the distributions can be ignored in comparison. Therefore, in summary, without any parallelization, for a sequence of $L$ tokens, HALC will cost approximately:
$L*T_{LVLM}+L*0.35*(T_{d}+n*T_{LVLM})=L*((1+0.35n)*T_{LVLM}+0.35T_{d})$ (29)
In practice, when $n=4$ and $T_{d}$ is relatively much smaller than $T_{LVLM}$ (the detection model Grounding DINO we used was based on the Swin-Tranformer [^8] https://huggingface.co/docs/transformers/model_doc/swin with 341M parameters), we expect HALC to cost around 2.4x of the normal greedy decoding time expense.
However, the decoding passes for the extra $n$ FOVs can essentially run in parallel as they do not depend on each other. With parallelization, the time cost with $n$ FOV decoding is equal to the time cost for 1 FOV decoding, so the expected time cost will be only approximately 1.35x of the greedy decoding. When the detection model time can not be ignored and in the worst case it is the same as the decoding step time (which is unlikely as the LVLMs we experimented with are 7B), the expected time cost would be 1.7x of the normal greedy decoding.
Appendix C Experimentation Details
C.1 Experimental Setups
@@
C.2 Empirical Studies on Optimal Visual Contexts
@@
Appendix D MME Experiment Details
The experiment details mostly follow Appendix C.2, where we adopt each sub-task of 30 images from the MME benchmark dataset [^15] https://github.com/BradyFU/Awesome-Multimodal-Large-Language-Models/tree/Evaluation, and reconstruct the question prompt following offline POPE. Specifically, instead of simply asking a question with a binary yes/no answer, we first ask the decoder to generate a detailed caption of the provided image and then check whether the target positive/negative word existes in the caption. The detailed results are reported in Table LABEL:tab:MME_result. The corresponding figure result is shown in Fig. 5.